OpenStudy (anonymous):

given the proportion a/b = 8/50 which ratio completes the equivalent proportion a/8=?

OpenStudy (anonymous):

1? if a/b is equal to 8 over 50 it would have to be the same fraction or a simplified one. but since they are referencing it to another equation the value must be known, so it would have to be 8 which would make it 8/8

OpenStudy (anonymous):

im not positive its just common knowledge

OpenStudy (anonymous):

no its not one.

OpenStudy (anonymous):

then i dont know my last guess would be 1.28 but other then that i dont have a clue :/

OpenStudy (anonymous):

the answer would be a fraction, not decimal.

OpenStudy (anonymous):

@phi @UnkleRhaukus

OpenStudy (nurali):

a/b = 8/50 a/8=b/50

OpenStudy (superdavesuper):

@Nurali has the right answer :)